OpenStudy (anonymous):

Simplify i^ 41 Answers: 1 -i -1 i

OpenStudy (anonymous):

i

OpenStudy (anonymous):

i

OpenStudy (anonymous):

reason is that it goes \[i\] \[i^2=-1\] \[i^3=-i\] \[i^4=1\] so repeats ever four. you have \[i^{41}\] so the only question is what is the remainder when you divide 41 by 4. the remainder is 1, so you know \[i^{41}=i^1=i\]